The product of two positive numbers is 48. The larger number is 13 more than the smaller. Find both numbers
Found 2 solutions by lwsshak3, Alan3354:
Answer by lwsshak3(11628)

The product of two positive numbers is 48. The larger number is 13 more than the smaller. Find both numbers
***
let x=smaller number
x+13=larger number
x(x+13)=48
x^2+13x=48
x^2+13x-48=0
(x+16)(x-3)=0
x=3
x+13=16
smaller number=3
larger number=16

Answer by Alan3354(69443)

The product of two positive numbers is 48. The larger number is 13 more than the smaller. Find both numbers
==============
Try pairs of factors of 48, find a pair that differ by 13.
---
1*48 NG
2*24 NG
4*12 NG
etc